OpenStudy (anonymous):

Can someone please teach me how to solve this? Solve the equation 4(-2x+1)=6-(2x-4)

OpenStudy (anonymous):

Just distribute, it will help a lot Ex: 4(-2x+1) = -8x+4 can you do the other for me?

OpenStudy (anonymous):

@tryhardhomework

OpenStudy (anonymous):

6-2x+4 right? It's probably wrong because that's what I've been doing

OpenStudy (anonymous):

strange... you would think that s how it would distributed

OpenStudy (anonymous):

Found the problem. I've been using negative 4 on the opposite site of the equation. Thank you though. I appreciate it.

OpenStudy (anonymous):

oh glad you figured it out :) what was your mistake again so i can learn from it? :)

OpenStudy (anonymous):

Well I coppied the equation from my book into here, but then I realized that I wrote down 4(-2x-4) instead of 4(-2x+1). I just had a minor brain lapse is all. Not sure if there's much to learn from that. :)

OpenStudy (anonymous):

ah okay thanks :) minor brain lapses seem to be common for me too

OpenStudy (anonymous):

You don't seem like it :)

OpenStudy (anonymous):

lolz everything is not always what it seems :P

OpenStudy (anonymous):

That's the Internet for you in one sentence haha

OpenStudy (anonymous):

ans is 1 for this equation

OpenStudy (anonymous):

@rajatgang please don't provide answers, thank you :)

OpenStudy (anonymous):

den what you want

OpenStudy (anonymous):

a process before the answer. Read the code of conduct, attached here: http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

oops sorry dude............its simple 4(-2x+1)=6-(2x-4) this is our question first of all we will multiply it 4*(-2x)+4*1 = 6-2x+4 -8x+4=-2x+10 now we will take like terms together i.e -8x+2x=10-4 -6x=-6 minus get cut 6x=6 x=1 hope u get it

OpenStudy (anonymous):

lol its not my question... but thanks for showing your process :) I am glad that you learned, make sure you always show a process, or else you may get in trouble. Cheers

OpenStudy (anonymous):

yup nd thnx .............but i think you also dnt knw the solution of this eq..nw u also knwing hw to solve these types of equations...nw u can also help others...... cheers

OpenStudy (anonymous):

Here is a step by step solution: http://www.symbolab.com/solutions/linear_equations?query=4(-2x%2B1)%3D6-(2x-4)

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."